Thank you for answering my question before, but my teacher said it was still wrong. The main word in the equation is 'hemisphere' she said....so now I am very confused =(. The question is: find the surface area of the hemisphere with radius 3 cm in terms of pi. it is not 36 pi cm 2.
Answer by stanbon(75887)

find the surface area of the hemisphere with radius 3 cm in terms of pi. it is not 36 pi cm 2.
----------------
The surface area of a sphere is 4(pi)r^2
-------------------------------------
Your Problem:
The surface area of a hemisphere = (1/2)4(pi)*3^2
---
= 2(pi)9
= 18(pi)cm^2
---
= 56.55.. cm^2
=======================
Cheers,
Stan H.